HETAV PATEL

Grade 11,

Two bottles of A and B contains 1M and 1m aqueous solutions (d=1g/ml) of sulphuric acid respectively. Compare the concentrations of bottle A and B. [AC ch5 pg 31 Q28]

Two bottles of A and B contains 1M and 1m aqueous solutions (d=1g/ml) of sulphuric acid respectively.
 
Compare the concentrations of bottle A and B.
 
 
[AC ch5 pg 31 Q28]

Grade:12th pass

1 Answers

Naveen Kumar
9 years ago
To compare convert both in same unit of concentratin (in molality).
{ solution=H2SO4+H2O}
In A, 1M means that 1mol of H2So4 is in 1000ml of solution.
So 1 mol H2SO4 is in 1000g of solution(as density is 1g/ml)
Now in 1000g of solution, mass of H2SO4=1 mol=98g
⇒1mol of H2SO4 is in (1000g-98g= 902g) of water.
Hence molality of H2SO4 in A=(1/902)*1000m=1.108m>molality in bottle B
So bottle A is more concentrated of H2So4 as compared to bottle B.